2. The City of Weston is preparing its budget for calendar year 2020. After estimating revenues from all other sources, the City calculates that it must raise $7,000,000 from property taxes. You are given the following information regarding the tax rate:

            Property taxes to be collected                                                              $7,000,000

            Estimated uncollectible property taxes                                                  $70,707

            Total assessed value of property at beginning of 2020                   $65,000,000

            Expected reduction in assessed value from appeals                           $200,000

             Assessed value of City property, not subject to tax                           $1,400,000

            Adjustments to assessed values for senior citizen exemptions        $1,000,000

Required:

  1. Compute the amount of property taxes required to be levied.
  2. Compute the tax rate per $100 of net assessed valuation.
  3. Determine the amount of property tax that a home owner whose property is assessed at $35,000 will have to pay.
Answer #1

Amount of Property Taxes to be levied are calculated as follows:

Property taxes to be collected = $7,000,000

Add: Estimated uncollectible property taxes = $70,707

Amount of property taxes required to be levied = $7,070,707

Calculation of Net Assessable Value :

Total assessed value of property at beginning of 2020 = $65,000,000

Less: Expected reduction in assessed value from appeals = $200,000

Less: Assessed value of City property, not subject to tax = $1,400,000

Less: Adjustments to assessed values for senior citizen exemptions = $1,000,000

Net Assessable Value = $62,400,000

Tax Rate per $100 of net assessed valuation = 7,070,707*100/62,400,000

= 11.3313% (approx.)

The amount of property tax that a home owner whose property is assessed at $35,000 will have to pay = 35,000*11.3313% = $3,966 (approx.)